OpenStudy (anonymous):

\[\sqrt[3]{24x^8}\] Simplify by factoring assume all expressions under radicals reprisent nonnegative number

OpenStudy (anonymous):

x^8 ?

OpenStudy (anonymous):

yes 8

OpenStudy (anonymous):

\[\sqrt[3]{2^3*3*x^3*x^3*x^2}=2x^2\sqrt[3]{3x^2}\]?

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

urwc